# Rapid near-inertial wave generation at topography in response to surface wind forcing

### v2

@@ Abstract
Internal waves propagate on the ocean stratification, carrying energy and redistributing momentum through the ocean interior. The two most significant sources of these waves in the ocean are surface winds and oscillatory tidal flow across topography. We propose a hybrid of these two mechanisms, in which wind-induced flows at the sea surface are rapidly communicated to the seafloor via the de-phasing of horizontally propagating vertical internal wave modes. In the presence of topography, the resulting oscillatory bottom velocity may then generate internal waves in a similar manner to the barotropic tide. We investigate this mechanism in an idealised numerical model of a storm passing over a mid ocean ridge, and perform several perturbation experiments in which ocean and wind properties are varied. Bottom-generated near-inertial internal waves are identified propagating away from the ridge in the wake of the storm. Perturbation experiments were performed when altering the shape of the topography and properties of the storm, with energy flux ranging from $0.01$ to $0.2~{}\mathrm{mW}~{}\mathrm{m}^{-2}$ or up to $10\%$ the magnitude of the energy flux from surface generated near-inertial waves in our domain. Our results suggest a need for further investigation in less idealised scenarios to more precisely quantify this novel mechanism of deep ocean wave generation, and how it may affect abyssal mixing. We argue that this mechanism could explain some of the unaccounted for bottom near-inertial wave signals found in ocean observations and numerical studies.
@@
The ocean is stratified: denser waters are in the deep ocean and lighter waters closer to the surface. How and where waters of different densities mix is important as this contributes to global ocean circulations and transport of heat from the tropics to the poles. This ocean mixing is influenced greatly by the winds and tides that churn up the ocean near boundaries and generate internal waves, the breaking of which is thought to be the major source of turbulent mixing in the ocean interior. These waves propagate in the interior of the ocean rather than just at the the surface, and are made possible by the ocean’s stratification. These waves are very difficult to observe and so our understanding of their generation, propagation and breaking is an active field of research. Here, we demonstrate a novel way that near-inertial waves are generated where storm systems blow above submerged features on the seafloor. We find that these waves could be up to $10\%$ as strong as surface generated near-inertial waves which are known to be a large contributor to mixing. We argue that this novel internal wave generation mechanism is worthy of further study to better understand how they might fit into the ocean mixing picture.
@@ 1 Introduction
Internal waves, which manifest as mixed horizontal and vertical oscillations in a rotating stratified fluid, permeate the world’s oceans. Internal waves are important to the fields of physical oceanography and climate science more broadly due to their significant contribution to ocean mixing and momentum transfer (Sarkar and Scotti, 2017; Thomas and Zhai, 2022; Musgrave et al., 2022; Alford et al., 2016; Bell, 1975). Global estimates of the energy contained by the internal wave field are supported by sparse in-situ observations and satellite altimetry (Whalen et al., 2018; Zhao et al., 2016). However, capturing the spatial variability of internal waves, and in turn their contribution to local episodic mixing, remains an ongoing challenge in oceanography, particularly in the deep ocean where observations are especially sparse. A further complication is that some types of internal wave can contribute both to the local mixing at the generation site, while others can carry their energy great distances to enhance mixing in remote locations (Waterhouse et al., 2014).
@@
The bulk of internal waves begin their life-cycle at the boundary of the ocean – either at the surface generated via winds (Thomas and Zhai, 2022), or near the bottom as a result of flow over topography (Musgrave et al., 2022). In the first case, inertial oscillations at the surface are readily excited by the wind, which in turn generate near-inertial waves that can propagate downwards, beyond the mixed layer (D’Asaro, 1985). These wind-driven, surface generated NIWs are responsible for between 0.3-1.1 TW of energy flux at the surface (Jiang et al., 2005; Rimac et al., 2013). The exact mechanism by which the near-inertial energy escapes the mixed layer depends significantly on the length scale of the wind forcing. Forcing by small, intense storm systems (Gill, 1984) generates low-modes which propagate thousands of kilometers horizontally away from their generation site (Alford, 2003). For larger scale winds, the near-inertial oscillations instead need to be downscaled by $\beta$-refraction or mesoscale interactions (Thomas and Zhai, 2022) in order to escape the mixed layer. These waves propagate at very low vertical group speeds and thus shallow angles, with observational studies suggesting that vertical propagation takes weeks to bring such waves to the seafloor after generation (Kunze and Sanford, 1984; Ma et al., 2022). Here, they may scatter, dissipate or reflect depending on the local topography (Igeta et al., 2009).
@@
In the second case, topography can also play a role in the generation of internal waves. When deep ocean flows interact with topography, this can lead to vertical motions, and thereby the radiation of internal waves due to the restoring effect of the stratification. Examples of such interactions include the barotropic tides generating internal tides (Bell, 1975), as well as lee waves excited by geostrophic currents (Nikurashin and Ferrari, 2010), which are thought to have globally-integrated energy fluxes of 0.9-1.5 TW and 0.05-0.85 TW respectively (Nikurashin and Ferrari, 2011; Egbert and Ray, 2000; Waterhouse et al., 2014; Bennetts et al., 2024).
@@
There is an indirect connection between wind-generated and topographically-generated internal waves. Wind forcing energises both long-lived and transient geostrophic currents (Wunsch, 1998), which in turn shed energy at the ocean bottom via topographic lee waves (Nikurashin and Ferrari, 2010). Nikurashin and Ferrari (2011) estimated that $20\%$ of energy input to geostrophic flows through synoptic scale wind forcing was converted to lee wave generation. Resonant interactions of the lee waves and geostrophic flows (Nikurashin and Ferrari, 2010) drive near-inertial energy and enhanced mixing near topographic features (Hu et al., 2020; Brearley et al., 2013; Liang and Thurnherr, 2012).
@@
Such bottom generated wave energy may help to explain the results of Waterhouse et al. (2022), in which global lowered acoustic Doppler current profiler observations identified large areas of the ocean where up-going waves (identified by the sign of rotary shear) exceeded down-going waves in the bottom 1000 m. As rotary shear is usually dominated by near-inertial internal waves, the results by Waterhouse et al. (2022) indicate the presence of bottom generated near-inertial waves in these areas. The broad extent of regions with enhanced upgoing waves in the bottom $1000~{}\mathrm{m}$ suggests there may be contributions from other as-yet unknown mechanisms for the generation of near-inertial waves at the seafloor (e.g., also see Alford et al., 2016). Further supporting this hypothesis are recent observations suggesting generation of near-inertial waves at the seafloor within hours of a storm passing overhead. van Haren (2020) found evidence for enhanced mixing in the depths of the Mariana Trench, with a mooring detecting patterns in the stratification similar to that of breaking internal waves within hours of the storm. The rapid timescale and close proximity of the mixing to the storm suggest that neither of the two cases discussed above — that is, the interaction with topography of lee wave-geostrophic flows or the downwards propagation of near-inertial waves initially generated in the mixed layer — are able to account for the observations. As such, a new mechanism is needed.
In this work, we investigate a process for near-inertial internal wave generation at the seafloor that can be thought of as a combination of two well studied generation mechanisms: near-inertial wave modes arising from an impulsive surface forcing (Gill, 1984) and internal tides arising from barotropic tidal flow over topography (Bell, 1975). We demonstrate that wind stress at the ocean’s surface, close to the inertial frequency, can almost instantaneously generate upwards-propagating internal waves at bottom topography near the storm, in addition to the more familiar near-inertial wave modes generated directly by wind forcing of the mixed layer.
@@
The mechanism for the rapid excitation of bottom velocities from a surface perturbation is that of de-phasing vertical modes. Gill (1984) studied the ocean response to a short-lived wind event, like the passage of a storm, that instantaneously accelerates the mixed layer and imprints onto multiple internal wave vertical modes. The rapid de-phasing of these vertical modes results in the rapid communication of the surface layer perturbation downwards below the thermocline. Given the full depth nature of these vertical modes, we extend Gill’s work to consider a rapid response at the sea floor.
@@
In the presence of topography, we anticipate that the deep oscillatory flow driven by de-phasing vertical modes will generate internal waves in a similar fashion to the barotropic tide (Bell, 1975). We expect these waves to be present immediately in the wake of a storm, and as such, they differ from the usual topographic scattering mechanism (discussed above) which is associated with a significant time lag due to wave propagation from surface to sea floor. Given that storms convert large amounts of wind energy into surface layer oscillations and downward propagating near-inertial waves (Sanford et al., 2011; Whalen et al., 2018), the coincidence of a storm directly above topographic features would provide the most pronounced generation. The hypothesised “wind-driven topographic near-inertial waves” (as we refer to them hereafter) have, to our knowledge, not been studied before and thus it is unknown whether they constitute a significant contribution to the ocean internal wave field and deep ocean mixing. To investigate this phenomenon, we run a 3D numerical ocean model with idealised wind forcing and bathymetry. Under various topographic and wind forcing scenarios, we calculate the energy flux associated with these waves, and compare the magnitude of energy flux to that of surface-generated near inertial waves excited directly by the wind.
@@
The next section reviews the theory of Gill (1984), and demonstrates the excitation of a bottom flow from the de-phasing mechanism. In section 3, we describe the model setup, as well as the method we used for separating the waves of interest from the rest of the flow. Section 4 includes a qualitative demonstration of these bottom near-inertial waves, followed by a quantitative analysis of the wave energy for various perturbation experiments. Finally, we conclude in section 5 by arguing that this mechanism could be responsible for some of the unexplained observations of excess bottom mixing and vertical shear.
@@ 2 Theory
Gill (1984) modelled the ocean’s immediate response to the passage of a storm as an initial velocity $u_{0}$ in the mixed layer, with zero velocity deeper in the water column. This initial condition thought of as an infinite set of orthonormal vertical modes each with a different vertical wavenumber $m_{n}$, which destructively interfere below the mixed layer. As these modes necessarily propagate at different speeds, they rapidly de-phase from one another, resulting in finite flows over the entire ocean depth.
To demonstrate this mechanism, consider an idealised 2D model of the ocean with a mixed layer of depth $h_{m}$, total depth $H$, and constant stratification $N$, just after the passage of a storm with characteristic horizontal length scale $L$ and associated wavenumber $k=2\pi/L$. The initial surface response to the storm when $t=0$ is given by an initial mixed-layer velocity with a maximum of $u_{0}$, while the interior is at rest; that is,
@@
Each of the vertical modes propagates at a different frequency and phase speed according to the dispersion relation for internal waves (Eq. (2)) meaning that lower modes propagate faster in the horizontal. Initially, the linear combination of modes with amplitudes prescribed by (4) sum to give $u_{0}$ in the mixed layer and zero at depth. But as time progresses, the different propagation speeds mean that the vertical modes de-phase from one another, and flow develops at depth.
Figure 1: a) Evolution of the initial surface layer velocity $u(x,z,t=0)$ for a reference case of $N=20|f|$ and a horizontal scale of $L=2\pi/k=100\;\mathrm{km}$. b) The bottom velocity response corresponding to the reference conditions, as well as for storms with decreased and increased horizontal scale $L$. Velocities have been non-dimensionalised against $u_{0}$
The manner in which the de-phasing happens depends on the differences in frequency $\omega_{n}$ and in turn on stratification $N$ and length scale $L$. Figure 1 shows the vertical structure of the solution at $x=0$ for $h_{m}=50\;\mathrm{m}$, $H=4000\;\mathrm{m}$, $|f|=10^{-4}\;\mathrm{s}^{-1}$ and $N=20|f|$ over 5 inertial periods. In this case, flow develops throughout the water column, including at the bottom, within 10% of an inertial period. Fig. 1b shows that varying $L$ affects both the amplitude and frequencies of resultant bottom flow, with the solution tending towards purely inertial oscillations as $\omega$ tends towards $f$. As propagating waves can be generated at topography with super-inertial oscillations, of particular interest are the stronger super-inertial motions that can be seen in the cases with smaller horizontal length scale.
The thickness of the surface layer $h_{m}$ plays a role in re-distributing the energy between vertical modes, but does not affect their fundamental de-phasing behaviour. This simple model demonstrates that a wind-induced surface perturbation in velocity — of the correct scale — can almost instantaneously generate super-inertial oscillatory flow at the sea floor. Existing theory for internal tides can now be applied in evaluating the likely impact of the wind-induced oscillatory flow on wave generation at the seafloor. The conversion rate $C$ of energy from an oscillatory flow of amplitude $U_{0}$ to buoyancy frequency $N$, topography height $h$, and horizontal wavenumber $k$, is given by (Jayne and St. Laurent, 2001):
@@
where $\rho_{0}$ is the reference density. Although topographically-generated near-inertial waves differ from internal tides in that the oscillatory flow excited by winds is not of constant amplitude or frequency, Eq. (5) holds for our study to leading order if $U_{0}$ refers instead to the characteristic amplitude of the wind-induced velocity. Our numerical results are evaluated against this scaling relation in section 5.

### v4

# Rapid topographic scatter of near-inertial waves generated by storms

@@ Abstract
Internal waves propagate on the ocean’s stratification, carrying energy and redistributing momentum through the ocean. When internal waves break, they contribute to diapycnal mixing in the ocean interior, but this breaking behaviour depends upon the scale of the waves. Low-mode internal waves have larger horizontal and vertical scales, and thus break less readily than higher-mode waves. The scattering of internal waves by topography is an important mechanism in transferring internal wave energy to smaller scales that are more conducive to wave breaking and mixing processes. In this study, we propose and investigate a mechanism in which storm-generated low-mode internal waves scatter at topography. We hypothesise that horizontally propagating internal wave modes generated by strong winds (i.e., due to a storm) can rapidly dephase; these dephased waves can then be scattered from topography, resulting in higher-mode upward-propagating waves within hours of the passage of a storm. We investigate this phenomenon in an idealised numerical model of a storm passing over a prominent ridge. Bottom-scattered near-inertial internal waves propagate away from the ridge rapidly in the wake of the storm. We perform several perturbation experiments varying the properties of the ocean, the winds and the topography. The bottom-scattered waves exhibit spatial downscaling, and have an energy flux equivalent to $10\%$ the magnitude of the energy flux from surface-generated near-inertial waves in our domain. Although small in a globally averaged sense, we argue that the topographic scatter of storm-generated near-inertial waves could account for the unexplained near-inertial wave signals found in ocean observations and numerical studies.
@@
The ocean is stratified: denser waters are in the deep ocean and lighter waters closer to the surface. How and where waters of different densities mix is important as this contributes to global ocean circulation and how the ocean transports heat from the tropics to the poles. This ocean mixing is influenced greatly by the winds and tides that churn up the ocean near boundaries and generate internal waves, the breaking of which is thought to be the major source of turbulent mixing in the ocean interior. These waves propagate in the interior of the ocean rather than just at the surface, and are made possible by the ocean’s stratification. These waves are very difficult to observe and so our understanding of their generation, propagation and breaking is an active field of research. Here, we demonstrate that internal waves generated by strong storms, like typhoons or cyclones, can rapidly scatter from nearby topography at the sea floor. Previous studies looked at waves generated by larger-scale winds rather than individual storms and suggested that the waves take weeks to reach the seafloor. Here, we show that storm-generated waves scatter at the seafloor within hours. Our study could help explain some of the bottom-intensified wave activity that has been observed within hours of the passage of a storm. We find that the waves propagating upwards from topography have smaller spatial scales, and up to $10\%$ the energy flux compared to the surface-generated waves that they scatter from. We argue that this rapid wave-scattering mechanism is worthy of further study to better understand how they might fit into the ocean mixing picture.
@@ 1 Introduction
Internal waves, which manifest as mixed horizontal and vertical oscillations in a rotating stratified fluid, permeate the world’s oceans. Internal waves are important to the fields of physical oceanography and climate science more broadly due to their significant contribution to ocean mixing and momentum transfer (Bell, 1975; Alford et al., 2016; Sarkar and Scotti, 2017; Thomas and Zhai, 2022; Musgrave et al., 2022). Global estimates of the energy contained by the internal wave field are supported by sparse in-situ observations and satellite altimetry (Zhao et al., 2016; Whalen et al., 2018). However, capturing the spatial variability of internal waves, and in turn their contribution to local episodic mixing, remains an ongoing challenge in oceanography, particularly in the deep ocean where observations are especially sparse. A further complication is that some types of internal wave can contribute to the local mixing at the generation site, while others can carry their energy great distances to enhance mixing in remote locations (Waterhouse et al., 2014).
@@
The bulk of internal waves begin their life-cycle at the boundary of the ocean – either at the surface, generated via winds (Thomas and Zhai, 2022), or near the bottom as a result of flow over topography (Musgrave et al., 2022). In the first case, inertial oscillations at the surface are readily excited by the wind, which in turn generate near-inertial waves that can propagate downwards, beyond the mixed layer (D’Asaro, 1985). These wind-driven, surface-generated near-inertial waves (NIWs) are responsible for between 0.3-1.1 TW of energy flux at the surface (Jiang et al., 2005; Rimac et al., 2013). The exact mechanism by which the near-inertial energy escapes the mixed layer depends significantly on the length scale of the wind forcing. Forcing by small, intense storm systems (Gill, 1984) generates low-modes which propagate thousands of kilometers horizontally away from their generation site (Alford, 2003). For larger scale winds, the near-inertial oscillations instead need to be downscaled by $\beta$-refraction or mesoscale interactions (Thomas and Zhai, 2022) in order to escape the mixed layer. These waves propagate at very low vertical group speeds and thus shallow angles, with observational studies suggesting that vertical propagation takes weeks to bring such waves to the seafloor after generation (Kunze and Sanford, 1984; Ma et al., 2022). Here, they may scatter, dissipate or reflect depending on the local topography (Igeta et al., 2009).
@@
In the second case, topography generates internal waves directly. When deep ocean flows interact with topography, this can lead to vertical motions, and thereby the radiation of internal waves due to the restoring effect of the stratification. Examples of such interactions include the barotropic tides generating internal tides (Bell, 1975), as well as lee waves excited by geostrophic currents (Nikurashin and Ferrari, 2010), which are thought to have globally-integrated energy fluxes of 0.9-1.5 TW and 0.05-0.85 TW respectively (Nikurashin and Ferrari, 2011; Egbert and Ray, 2000; Waterhouse et al., 2014; Bennetts et al., 2024).
@@
There is an indirect connection between wind-generated and topographically-generated internal waves. Wind forcing energises both long-lived and transient geostrophic currents (Wunsch, 1998), which in turn shed energy at the ocean bottom via topographic lee waves (Nikurashin and Ferrari, 2010). Nikurashin and Ferrari (2011) estimated that $20\%$ of energy input to geostrophic flows through synoptic scale wind forcing was converted to lee wave generation. Resonant interactions of the lee waves and geostrophic flows (Nikurashin and Ferrari, 2010) drive near-inertial energy and enhanced mixing near topographic features (Hu et al., 2020; Brearley et al., 2013; Liang and Thurnherr, 2012). Given that the wind-driven geostrophic currents responsible for lee wave generation vary on timescales of months and years, this mechanism represents a persistent, steady source of internal wave generation.
@@
In contrast, recent observations have found rapid generation of near-inertial waves at the seafloor within hours of a storm passing overhead. van Haren (2020) found evidence for enhanced mixing in the depths of the Mariana Trench, with a mooring detecting patterns in the stratification similar to that of breaking internal waves within hours of a category 4 typhoon. The rapid timescale and close proximity of the mixing to the storm suggest that neither of the two cases discussed above — that is, the interaction of lee waves with geostrophic flows nor the downwards propagation of near-inertial waves with low vertical group speeds excited by large scale winds — are able to account for the observations. As such, a new mechanism is needed.
In this work, we investigate the rapid scattering of near-inertial internal waves at the seafloor in the wake of a storm system. We hypothesise that the sharp horizontal gradients of storm systems allow for the generation of full-depth vertical modes, unlike other near-inertial waves generated by larger scale wind patterns as discussed in Kunze and Sanford (1984) and Ma et al. (2022). Other full-depth internal wave modes, like the baroclinic tide, are known to scatter at topography, which can redistribute their energy to smaller horizontal scales (Peter Müller and Naihuai Xu, 1992; Laurent and Garrett, 2002; Bühler and Holmes-Cerfon, 2011; Kelly et al., 2013). Like these other full-depth internal waves, we demonstrate that the near-inertial waves generated by a storm can also scatter from local topography, with upwards propagating internal waves appearing almost immediately. This short timescale of the bottom response is of particular interest, as it makes this mechanism a contender for explaining the aforementioned enhanced bottom mixing coinciding with the typhoon.
@@
The mechanism for this rapid topographic scattering is the dephasing of vertical modes. Gill (1984) studied the ocean response to a short-lived wind event, like the passage of a storm, that instantaneously accelerates the mixed layer and imprints onto multiple internal wave vertical modes. The dephasing of these vertical modes results in the rapid communication of the surface layer perturbation downwards below the thermocline. Given the full depth nature of the vertical modes generated by storm systems, we extend Gill’s work to consider a rapid response at the sea floor.
In the presence of topography, we anticipate that the deep oscillatory motions driven by dephasing vertical modes will generate internal waves in a similar fashion to internal tides (Bell, 1975). Like internal tides, this generation process can equally be thought of as the scattering of a large scale wave (barotropic tide, or surface-generated NIW) to a smaller scale wave (baroclinic tide, or topographically scattered NIW).
As discussed above, the scattering of near-inertial waves generated by large-scale winds is associated with a significant time lag due to them not exciting full-depth modes, and the low propagation angle of the waves as they escape the mixed layer. In contrast, the smaller horizontal scales of storm-generated near-inertial waves, which are the focus of this study, readily excite full-depth modes, and in turn we expect a rapid bottom response as these modes dephase. In the presence of sufficiently narrow and prominent topographic features, we further expect the waves to scatter to higher modes, facilitating energy downscaling which may ultimately lead to additional bottom mixing.
@@
Given that storms convert large amounts of wind energy into surface layer oscillations and downward propagating near-inertial waves (Sanford et al., 2011; Whalen et al., 2018), the coincidence of a storm directly above topographic features would provide the most pronounced generation. The hypothesised topographic scatter of storm-generated near-inertial waves have, to our knowledge, not been studied before and thus it is unknown whether they constitute a significant contribution to the ocean internal wave field and deep ocean mixing. To investigate this phenomenon, we run a numerical ocean model with idealised wind forcing and bathymetry. Under various topographic and wind forcing scenarios, we calculate the energy flux associated with these waves, and compare the magnitude of energy flux to that of surface-generated near inertial waves excited directly by the wind.
@@
The next section reviews the theory of Gill (1984), and demonstrates the excitation of a bottom flow from the dephasing mechanism. In section 3, we describe the model setup, as well as the method we used for separating the waves of interest from the rest of the flow. Section 4 includes a qualitative demonstration of these bottom scattered waves, followed by a quantitative analysis of the wave energy for various perturbation experiments. Finally, we conclude in section 5 by arguing that this mechanism could be responsible for some of the unexplained observations of excess bottom mixing and vertical shear.
@@ 2 Theory
Gill (1984) modelled the ocean’s immediate response to the passage of a storm as an initial velocity $u_{0}$ in the mixed layer, with zero velocity deeper in the water column. This initial condition can be thought of as an infinite set of orthonormal vertical modes each with a different vertical wavenumber $m_{n}$, which destructively interfere below the mixed layer. As these modes necessarily propagate at different speeds, they rapidly de-phase from one another, resulting in finite flows over the entire ocean depth.
To demonstrate this mechanism, consider an idealised hydrostatic 2D model of the ocean with a mixed layer of depth $h_{m}$, total depth $H$, and constant stratification $N$, just after the passage of a storm with characteristic horizontal length scale $L$ and associated wavenumber $k=2\pi/L$. The initial surface response to the storm when $t=0$ is given by an initial mixed-layer velocity with a maximum of $u_{0}$, while the interior is at rest; that is,
@@
Each of the vertical modes propagates at a different frequency and phase speed according to the dispersion relation for internal waves (2) meaning that lower modes propagate faster in the horizontal. Initially, the linear combination of modes with amplitudes prescribed by (4) sum to give $u_{0}$ in the mixed layer and zero at depth. But as time progresses, the different propagation speeds mean that the vertical modes de-phase from one another, and flow develops at depth.
Figure 1: a) Evolution of the initial surface layer velocity $u(x,z,t=0)$ for a reference case of $N=20|f|$ and a horizontal scale of $L=2\pi/k=100\;\mathrm{km}$. b) The bottom velocity response corresponding to the reference conditions, as well as for storms with decreased ($20~{}\mathrm{km}$) and increased ($1000~{}\mathrm{km}$) horizontal scale $L$. Velocities have been non-dimensionalized against $u_{0}$, which in this case is $0.1\;\mathrm{m}\;\mathrm{s}^{-1}$ over a $200\;\mathrm{m}$ thick mixed layer $h_{m}$.
The manner in which the dephasing happens depends on the differences in frequency $\omega_{n}$ and in turn on stratification $N$ and length scale $L$. Figure 1 shows the vertical structure of the solution at $x=0$ for $h_{m}=200\;\mathrm{m}$, $H=4000\;\mathrm{m}$, $|f|=10^{-4}\;\mathrm{s}^{-1}$ and $N=20|f|$ over 5 inertial periods. In this case, flow develops throughout the water column, including at the bottom, within 10% of an inertial period. Figure 1b shows that varying $L$ affects both the amplitude and frequencies of resultant bottom flow, with the solution tending towards purely inertial oscillations as $L$ increases. As propagating waves can be generated at topography with super-inertial oscillations, of particular interest are the stronger super-inertial motions that can be seen in the cases with smaller horizontal length scale.
The thickness of the surface layer $h_{m}$ plays a role in re-distributing the energy between vertical modes, but does not affect their fundamental dephasing behaviour. This simple model demonstrates that a wind-induced surface perturbation in velocity — of the correct scale — can almost instantaneously generate super-inertial oscillatory flow at the sea floor. Existing theory for internal tides can now be applied in evaluating the likely impact of the wind-induced oscillatory flow on secondary wave generation — or scatter — at the seafloor. The conversion rate $C$ of energy from an oscillatory flow of amplitude $U_{0}$ to buoyancy frequency $N$, topography height $h$, and horizontal wavenumber $k$, based on linear wave theory, is given by (Jayne and St. Laurent, 2001):
@@
where $\rho_{0}$ is the reference density. Although this mechanism differs from internal tides in that the oscillatory flow excited by winds is not of constant amplitude or frequency, (5) holds for our study to leading order if $U_{0}$ refers instead to the characteristic amplitude of the wind-induced velocity. Our numerical results are evaluated against this scaling relation in section 5.